The scores of Miss San Quintin candidates from seven judges were recorded as follows: 8.12, 9.10, 8.42, 9.14, 8.75, 9.17, and 8. 37

For numbers 4 – 6, refer to the given data above.

4. Find the 50th percentile or P50 of the judges’ scores.

M. 9.14 A. 8.75 T. 8.42 H. 8.12

5. What is the P30 of the judges’ scores?

M. 9.10 A. 8.75 T. 8.42 H. 8.37

6. Determine the 70th percentile of the judges’ scores.

M. 9.14 A. 9.10 T. 8.42 H. 8.12



1
Expert's answer
2022-05-09T11:10:41-0400

4. To get the percentile in a given set of data, the data should be sorted from lowest to highest. 


"8.12, 8. 37, 8.42, 8.75, 9.10, 9.14, 9.17"

"median=8.75"

A. 8.75


5. To get the percentile in a given set of data, the data should be sorted from lowest to highest. 


"8.12, 8. 37, 8.42, 8.75, 9.10, 9.14, 9.17"

"n=7"

"P30=\dfrac{30}{100}(7+1)=2.4"

We rounded down P30 to 2 and what it means is that the P30 is the 2nd rank of the sorted data set.

H. 8.37


6. To get the percentile in a given set of data, the data should be sorted from lowest to highest. 


"8.12, 8. 37, 8.42, 8.75, 9.10, 9.14, 9.17"

"n=7"

"P70=\dfrac{70}{100}(7+1)=5.6"

We rounded down P70 to 6 and what it means is that the P70 is the 6th rank of the sorted data set.

The 6th rank of the sorted data set is 9.14

M. 9.14